# Social media polarization during conflict: Insights from an ideological stance dataset on Israel-Palestine Reddit comments

### v1

@@ Introduction
  • Truncation: To manage computational efficiency and memory usage, texts were truncated to a maximum length of 700 characters, aligning with recommendations for handling lengthy sequences in transformer-based models [25]. Truncated texts were used in all models.
Dataset Split
@@
Neural Network Architectures
Deep learning models, including the RNN, LSTM [26], Bi-LSTM [27], and GRU [28], were trained and evaluated for the task. Among the various vectorization techniques explored, Bag of Words (BoW) has emerged as the most effective for these architectures. To enhance model performance, hyperparameter tuning was conducted, focusing on optimizing the configurations of the neural network architectures. Table 1 lists the tested hyperparameters and their respective ranges.
@@
Pre-trained Language Models
@@
  • BERT (Cased and Uncased): BERT is a transformer-based model which is pre-trained using a masked language model objective in order to capture bidirectional context in text. The cased version preserves case sensitivity, whereas the uncased version ignores it [25].
@@
  • XLM-RoBERTa: XLM-RoBERTa extends RoBERTa to cross-lingual tasks by leveraging massive multilingual datasets that allow it to achieve enhanced performance on multilingual tasks ranging across 100 languages [29].
@@
  • DistilBERT (Cased and Uncased): This is the distilled version of BERT which offers faster processing speeds and reduced model size while retaining 97 % of BERT’s performance [30].
@@
  • ELECTRA (Small and Base): This model employs a unique token-replacement detection mechanism for efficient pre-training which enables the model to train faster and learn from smaller datasets [31].
Large Language Models
@@
In politically sensitive datasets such as ours, nuanced language patterns and implicit stances often go unnoticed by traditional models. However, Large Language Models (LLMs) offer a promising solution, as they are pre-trained on extensive and diverse datasets, enabling them to generalize effectively to new contexts with minimal task-specific fine-tuning. Their ability to handle zero-shot, one-shot, and few-shot scenarios makes them particularly advantageous in low-resource settings where labeled data are limited. Additionally, LLMs are well-suited for processing longer and more complex narratives, which is especially relevant for our dataset comprising detailed Reddit comments with subtle political leanings. In this study, we conducted experiments using four Large Language Models, all accessed via the Huggingface platform. [^2] Huggingface platform: https://huggingface.co
@@
  • Mistral: Mistral is a dense and efficient transformer-based language model with 7 billion parameters. It is trained on a mixture of high-quality datasets using techniques such as group-query attention and multi-query key-value caching for efficiency. It demonstrates advanced reasoning and comprehension capabilities, often outperforming Llama 2 13B on standard benchmarks [32].
@@
  • Mixtral: Mixtral 8x7B is a Sparse Mixture of Experts (SMoE) model based on Mistral 7B which uses 8 experts per layer while activating only 13B parameters at a time [33].
@@
  • Gemma: Gemma is a family of lightweight open models based on the Gemini framework that incorporates multi-modal and multi-task training [34]. It offers two parameter settings- 2 and 7 billions. We used the 7B parameter version in our study.
@@
  • Falcon: Falcon is a series of causal decoder-only models with 7B, 40B, and 180B parameters, trained on over 3.5 trillion tokens of high-quality web data [35]. The 7B version was utilized in our study.
@@
We applied zero-shot, one-shot, three-shot, and five-shot prompting strategies [36] to the truncated text column of the dataset to identify the underlying ideological stance in each entry. These strategies evaluated the models’ abilities to handle varying levels of contextual information and infer classifications for the text.
@@
    Fig 3: Zero and One Shot Prompting Demonstration
@@
    Fig 4: Three and Five Shots Prompting Demonstration
Novel Prompting Strategies
This section explores the impact of various novel prompting techniques on the model performance. We used five novel prompting strategies, consisting of Re-read Zero-Shot and One-Shot, Meta-Prompting, Scoring & Reflective Re-read and Context Extraction. We evaluated these strategies only with Mixtral because its zero- and one-shot performance yielded the best results.
  • Re2 or Re-read: Re2 or re-read prompting strategy is designed to make the model check its classification and the text itself twice to reassess its initial output. This approach involves careful analysis and minimizes the likelihood of overlooking subtle contextual cues in its initial classification. We extended this approach to a one-shot setting where it reassesses its classification after being presented with an example during the initial classification. This addition combines the iterative strength of Re-read with the guidance provided in a one-shot setting. Figures 5 and 6 demonstrate the application of the Re-read prompting strategy in our context.
    Fig 5: Re-read Prompting Demonstration
    Fig 6: Re-read-One Shot Prompting Demonstration
@@
    Fig 7: Meta-Prompting (Self-Critique) Demonstration
@@
    Fig 8: Context Extraction Demonstration
@@
    Fig 9: Scoring and Reflective Re-read Demonstration
Experimental Setup
Hardware Details
The experiments were conducted on the Kaggle and Google Colab platform utilizing the free NVIDIA Tesla P100 and A100 GPU. This facilitated the training and evaluation of the models, particularly in handling computationally expensive processing.
Evaluation Metrics
To evaluate the performance of the models implemented for sentiment classification, we employed four widely recognized metrics: Accuracy, Precision (Macro), Recall (Macro), and F1 Score (Macro). These metrics were chosen to ensure a comprehensive analysis, particularly given the multi-class nature of the task.
Accuracy: Accuracy refers to the proportion of correctly predicted instances among the total instances evaluated and serves as an overall indicator of a model’s performance [37].
$\text{Accuracy}=\frac{TP+TN}{TP+TN+FP+FN}$ (1)
Where
  • TP: True Positives
  • TN: True Negatives
  • FP: False Positives
  • FN: False Negatives
In this study, accuracy provides a baseline understanding of model performance across the three stance categories [38].
Precision (macro): Precision evaluates the proportion of correctly predicted instances of a specific class relative to the total predicted instances for that class. Macro-averaging aggregates precision scores across all classes by assigning equal weights to each class, regardless of the number of instances [39, 40].
Macro $\text{Macro Precision}=\frac{1}{N}\sum_{i=1}^{N}\frac{TP_{i}}{TP_{i}+FP_{i}}$ (2)
Where
  • N: Total number of classes
  • $TP_{i}$: True positives for the i-th class
  • $FP_{i}$: False positives for the i-th class
Precision (Macro) ensures that smaller classes, such as potentially underrepresented sentiments, are given equal importance during the evaluation [38].
Recall (macro): Recall, or sensitivity, measures the proportion of correctly identified instances of a class among all actual instances for that class [39, 40].
Macro $\text{Macro Recall}=\frac{1}{N}\sum_{i=1}^{N}\frac{TP_{i}}{TP_{i}+FN_{i}}$ (3)
Where
  • $FN_{i}$: False negatives for the i-th class
Recall (Macro) is particularly critical in this study to evaluate the model’s ability to identify each sentiment class comprehensively, especially when class distributions are uneven [38].
F1 Score (macro): The F1 Score combines precision and recall into a single metric by calculating the harmonic mean. Macro F1 averages the F1 scores for all classes [41].
[…]
Where
  • $Precision_{i}$: Precision for the i-th class
  • $Recall_{i}$: Recall for the i-th class
Recall (Macro) is particularly critical in this study to evaluate the model’s ability to identify each sentiment class comprehensively, especially when class distributions are uneven [38].
Performance Evaluation
Baseline Models
Among the LLMs, only Mixtral 8x7B and Gemma 7B crossed the 60% accuracy threshold. The Mixtral 8x7B Zero Shot had higher accuracy, macro recall, and F1-score, while the Mixtral 8x7B One Shot had the highest macro precision.
@@
Category Model Test Accuracy Test Precision (macro) Test Recall (macro) Test F1 Score (macro)
@@
Large Language Models Mixtral 8x7B Zero Shot 0.6830 0.6612 0.6650 0.6628
@@
Table 3 lists the precision, recall, and F1 score of the best-performing Mixtral 7B Zero Shot. The model was observed to make a significant number of errors in classifying Neutral texts. It showed the best performance in classifying Pro-Israel texts, although the Pro-Palestine classifications were also substantially better than Neutral classifications.
Table 3: Class-wise performance of Mixtral 7B Zero Shot.
@@
Considering the performance scores of Mixtral 8x7B, we tested some prompt engineering techniques on the model to improve its base performance.
Novel Methods
Table 4 presents the results of the novel prompt engineering methods used. Among them, Scoring & Reflective Re-read achieved the highest performance with an accuracy of 74.08%, macro precision of 73.25%, macro recall of 73.09%, and a macro F1 score of 72. 93%. Although the one-shot re-read had an accuracy close to 70%, it fell short of the Scoring & Reflective Re-read across all metrics.
Table 4: Performance of Mixtral 7B with different prompt engineering techniques.
@@
Mixtral 8x7B - Re2 0.6374 0.7246 0.6634 0.6432
Mixtral 8x7B - Re2 (One-Shot) 0.6965 0.7144 0.6880 0.6896
@@
Mixtral 8x7B (Context Extraction) 0.6094 0.5969 0.6000 0.5971
Table 5 lists the class-wise performance of the Scoring & Reflective Re-read method. We can observe that the classification performance has increased significantly for all classes compared to the base zero-shot model. The most significant improvements were observed in classifying the neutral classes, where the most error occurred previously. These improvements suggest that the Scoring and Reflective Re-read achieves the best possible outcome out of all tested methods, thus making it our proposed method based on those experimental results.
@@
Conclusion
@@
This study provides a comprehensive analysis of ideological stance detection in social media data in the context of the Israel-Palestine conflict. By leveraging a unique manually annotated dataset of Reddit comments, we compared the performances of neural networks, pre-trained language models, and large language models (LLMs) using various evaluation metrics. The findings reveal the nuanced nature of ideological sentiment in conflict zones and underscore the importance of advanced models and prompt engineering techniques to achieve higher classification accuracy.
Among the models tested, Mixtral 8x7B with Scoring and Reflective Re-read prompting strategies demonstrated superior performance, achieving significant improvements in accuracy, precision, recall, and F1-scores across all stance categories. This research highlights the limitations of traditional models in capturing subtle and often implicit stances in politically sensitive datasets, and demonstrates the potential of novel prompting techniques to enhance classification outcomes.
@@
Future work can build upon this foundation by exploring multilingual datasets, integrating cross-cultural stances, and evaluating the ethical implications of ideological stance detection in highly polarized scenarios. The proposed methodologies and findings contribute to the broader discourse on the role of machine learning and natural language processing in understanding and addressing social media polarization in conflict zones.
Author Contributions
Conceptualization: Hasin Jawad Ali.
Methodology: Hasin Jawad Ali, S.M. Hozaifa Hossain.
Data Curation: Hasin Jawad Ali, S.M. Hozaifa Hossain, Ajwad Abrar.
Formal analysis: Hasin Jawad Ali, S.M. Hozaifa Hossain.
Writing – Original Draft Preparation: Hasin Jawad Ali, S.M. Hozaifa Hossain.
Writing – Review & Editing: Ajwad Abrar.
Visualization: Hasin Jawad Ali.
Validation: Ajwad Abrar, M. Firoz Mridha.
Supervision: Ajwad Abrar, M. Firoz Mridha.
Project Administration: Ajwad Abrar, M. Firoz Mridha.
Investigation: M. Firoz Mridha.

### v2

@@ 3.5 Data Processing
  • Truncation: To manage computational efficiency and memory usage, texts were truncated to a maximum length of 700 characters, aligning with recommendations for handling lengthy sequences in transformer-based models [devlin2019bert]. Truncated texts were used in all models.
3.6 Dataset Split
@@ 4 Methodology
4.1 Neural Network Architectures
Deep learning models, including the RNN, LSTM [Sherstinsky_2020], Bi-LSTM [cui2019deepbidirectionalunidirectionallstm], and GRU [dey2017gru], were trained and evaluated for the task. Among the various vectorization techniques explored, Bag of Words (BoW) has emerged as the most effective for these architectures. To enhance model performance, hyperparameter tuning was conducted, focusing on optimizing the configurations of the neural network architectures. Table 3 lists the tested hyperparameters and their respective ranges.
@@ 4.1 Neural Network Architectures
4.2 Pre-trained Language Models
@@
  • BERT (Cased and Uncased): BERT is a transformer-based model which is pre-trained using a masked language model objective in order to capture bidirectional context in text. The cased version preserves case sensitivity, whereas the uncased version ignores it [devlin2019bert].
@@
  • XLM-RoBERTa: XLM-RoBERTa extends RoBERTa to cross-lingual tasks by leveraging massive multilingual datasets that allow it to achieve enhanced performance on multilingual tasks ranging across 100 languages [conneau2020xlmr].
@@
  • DistilBERT (Cased and Uncased): This is the distilled version of BERT which offers faster processing speeds and reduced model size while retaining 97 % of BERT’s performance [sanh2019distilbert].
@@
  • ELECTRA (Small and Base): This model employs a unique token-replacement detection mechanism for efficient pre-training which enables the model to train faster and learn from smaller datasets [clark2020electra].
4.3 Large Language Models
@@
In politically sensitive datasets such as ours, nuanced language patterns and implicit stances often go unnoticed by traditional models. However, large language models (LLMs) offer a promising solution, as they are pre-trained on extensive and diverse datasets, enabling them to generalize effectively to new contexts with minimal task-specific fine-tuning. Their ability to handle zero-shot, one-shot, and few-shot scenarios makes them particularly advantageous in low-resource settings where labeled data are limited. Additionally, LLMs are well-suited for processing longer and more complex narratives, which is especially relevant for our dataset comprising detailed Reddit comments with subtle political leanings. In this study, we conducted experiments using seven LLMs, all accessed via the Huggingface platform. [^2] Huggingface platform: https://huggingface.co
@@
  • Mistral: Mistral is a dense and efficient transformer-based language model with 7 billion parameters. It is trained on a mixture of high-quality datasets using techniques such as group-query attention and multi-query key-value caching for efficiency. It demonstrates advanced reasoning and comprehension capabilities, often outperforming Llama 2 13B on standard benchmarks [jiang2023mistral7b].
@@
  • Mixtral: Mixtral 8x7B is a Sparse Mixture of Experts (SMoE) model based on Mistral 7B which uses 8 experts per layer while activating only 13B parameters at a time [jiang2024mixtralexperts].
@@
  • Gemma: Gemma is a family of lightweight open models based on the Gemini framework that incorporates multi-modal and multi-task training [gemmateam2024gemmaopenmodelsbased]. It offers multiple versions including Gemma 2B, Gemma 7B, Gemma 2-9B, Gemma 2-27B, Gemma 3-4B etc. We used Gemma 7B and Gemma 3-4B in our study which contain 8.54 and 4 billion parameters respectively.
@@
  • Falcon: Falcon is a series of causal decoder-only models that includes many versions including Falcon 7B, Falcon 40B, Falcon 180B, Falcon 2-11B, Falcon 3-3B etc., trained on over 3.5 trillion tokens of high-quality web data [Almazrouei2023TheFS]. We utilized Falcon 7B and Falcon 3-3B in our study.
  • Qwen: The Qwen series, developed by Alibaba Group, comprises Transformer-based large language models trained on diverse multilingual corpora encompassing web texts, books, and code [bai2023qwentechnicalreport]. There are many types of Qwen models including QWEN-CHAT, CODE-QWEN, CODE-QWEN-CHAT, and MATH-QWEN-CHAT. They also offer many parameter choices including 1,7B, 4B, 122B etc. We utilized Qwen 3-4B in our study.
@@
We applied zero-shot, one-shot, three-shot, and five-shot prompting strategies [brown2020languagemodelsfewshotlearners] to the truncated text column of the dataset to identify the underlying ideological stance in each entry. These strategies evaluated the models’ abilities to handle varying levels of contextual information and infer classifications for the text.
@@
    Figure 7: Zero and One Shot Prompting Demonstration
@@
    Figure 8: Three and Five Shots Prompting Demonstration
Novel Prompting and Classification Strategies
We used five novel prompting strategies, consisting of Re-read Zero-Shot and One-Shot, Meta-Prompting, Scoring & Reflective Re-read, and Context Extraction. These methods have not been previously used in this domain. We evaluated these strategies only with Mixtral since its zero- and one-shot performance yielded the best results.
  • Re2 or Re-read: Re2 or re-read prompting strategy is designed to make the model check its classification and the text itself twice to reassess its initial output. This approach involves careful analysis and minimizes the likelihood of overlooking subtle contextual cues in its initial classification. We extended this approach to a one-shot setting where it reassesses its classification after being presented with an example during the initial classification. This addition combines the iterative strength of Re-read with the guidance provided in a one-shot setting. Figures 9 and 10 demonstrate the application of the Re-read prompting strategy in our context.
    Figure 9: Re-read Prompting Demonstration
    Figure 10: Re-read-One Shot Prompting Demonstration
@@
    Figure 11: Meta-Prompting (Self-Critique) Demonstration
@@
    Figure 12: Context Extraction Demonstration
@@
    Figure 13: Scoring and Reflective Re-read Demonstration
5 Experimental Setup
5.1 Hardware Details
The experiments were conducted on the Kaggle and Google Colab platform utilizing the free NVIDIA Tesla P100 and A100 GPU. This facilitated the training and evaluation of the neural network and pre-trained transformer models, particularly in handling computationally expensive processing. Most LLMs, due to their massive sizes, were accessed via the Hugging Face Inference API and OpenRouter API. However, lighter LLMs including Qwen 3-4B, and Falcon 3-3B were downloaded and run locally.
5.2 Evaluation Metrics
To evaluate the performance of the models implemented for sentiment classification, we employed four widely recognized metrics: Accuracy [agustian2024new], Precision (Macro), Recall (Macro) [cahya2024improving, shakith2024enhancing], and F1 Score (Macro) [berger2020threshold]. Macro metrics were chosen to ensure a comprehensive analysis, particularly due to the uneven class distribution in the dataset [hote2023open]. On the other hand, accuracy provides a baseline understanding of model performance across the three stance categories. Detailed descriptions of evaluations metrics are mentioned in Appendix A.
@@ 5.2 Evaluation Metrics
Category Model Accuracy Precision Recall F1 Score
(Macro) (Macro) (Macro)
@@
Large Language Models Qwen3-4B Zero Shot 0.6583 0.6831 0.6110 0.6250
Qwen3-4B One Shot 0.6033 0.6098 0.6054 0.6033
Qwen3-4B Three Shot 0.5600 0.5614 0.5334 0.5391
Qwen3-4B Five Shot 0.5500 0.5618 0.5206 0.5186
Mixtral 8x7B Zero Shot 0.6830 0.6612 0.6650 0.6628
@@
Gemma3-4B Zero Shot 0.6583 0.6831 0.6110 0.6250
Gemma3-4B One Shot 0.6017 0.6262 0.5794 0.5774
Gemma3-4B Three Shot 0.6733 0.7086 0.6221 0.6230
Gemma3-4B Five Shot 0.6867 0.7085 0.6359 0.6437
@@
Falcon3-3B Zero Shot 0.4900 0.4914 0.4866 0.4886
Falcon3-3B One Shot 0.5183 0.5298 0.5476 0.5199
Falcon3-3B Three Shot 0.4700 0.5168 0.4646 0.4545
Falcon3-3B Five Shot 0.4850 0.5245 0.5030 0.4839
6 Performance Evaluation
6.1 Baseline Models
Among the LLMs, Mixtral 8x7B, Qwen3-4B, Gemma 7B, and Gemma 3-4B crossed the 60% accuracy threshold. Gemma 3-4B Five Shot had the highest accuracy, while Mixtral 8x7B Zero and One Shot had the highest macro recall, precision and F1-score. Considering the class imbalance in the dataset, and a comparatively smaller accuracy gain of Gemma 3-4B over Mixtral 8x7B, Mixtral 8x7B is selected as the best performing model for its superiority across macro metrics.
Table 5 lists the precision, recall, and F1 score of the best-performing Mixtral 7B Zero Shot. Figure 14 shows the confusion matrix. The model was observed to make a significant number of errors in classifying Neutral texts. It showed the best performance in classifying Pro-Israel texts, although the Pro-Palestine classifications were also substantially better than Neutral classifications.
Figure 14: Confusion Matrix of Mixtral 8x7B Zero Shot
Table 5: Class-wise performance of Mixtral 8x7B Zero Shot.
@@ 6.1 Baseline Models
Due to the superiority of the best performing model, we tested the novel prompt engineering techniques on the model to improve its base performance in the next section. We also show the result of LLM majority voting classification in the section after that.
6.2 Novel Methods
Table 6 presents the results of the proposed prompting strategies along with the LLM majority voting ensemble for comparison. Among them, Scoring & Reflective Re-read achieved the highest performance in accuracy with a score of 74.08%, in macro recall with a score of 73.09%, and in macro F1 with a score of 72.93%. However, LLM majority voting had a slightly higher macro precision with a score of 73.52%. Although the one-shot re-read had an accuracy close to 70%, it fell short of the Scoring & Reflective Re-read across all metrics. Considering superiority across all metrics and the difference in performance among the methods, Scoring & Reflective Re-read emerges as the best performing method.
Table 6: Performance of Proposed Prompting Strategies and the LLM Majority Voting Ensemble
@@ 6.2 Novel Methods
Mixtral 8x7B Re2 0.6374 0.7246 0.6634 0.6432
Mixtral 8x7B Re2 (One-Shot) 0.6965 0.7144 0.6880 0.6896
@@
Mixtral 8x7B Context Extraction 0.6094 0.5969 0.6000 0.5971
LLM Majority Voting 0.6267 0.7352 0.5886 0.6311
Figure 15: Confusion Matrix (Scoring+Re2)
Table 7 lists the class-wise performance of the Scoring & Reflective Re-read method and Figure 15 shows the confusion matrix. We can observe that the classification performance has increased significantly for all classes compared to the base zero-shot model. The most significant improvements were observed in classifying the neutral classes, where the most error occurred previously. These improvements suggest that the Scoring and Reflective Re-read achieves the best possible outcome out of all tested methods, thus making it our proposed method based on those experimental results.
@@
6.3 LLM Ensemble via Majority Voting
To examine whether combining predictions from multiple large language models improves classification robustness, we implemented an ensemble strategy based on majority voting. Specifically, predictions from three lightweight LLMs (Gemma 3-4B, Falcon 3-3B, and Qwen 3-4B) were aggregated for each test instance. These models were selected because they can be executed locally and represent diverse architectures among the evaluated LLMs.
For each comment, the stance predictions generated by the three models were collected, and the final label was determined based on the majority class among the predictions. In cases where no majority occurred, the prediction of the best-performing individual model was selected.
The performance of the ensemble model is reported in Table 6. The majority voting approach achieved an accuracy of 62.67%, a macro precision of 73.52%, a macro recall of 58.86%, and a macro F1-score of 63.11%. While this ensemble strategy slightly improved macro precision compared to several individual LLM configurations, it did not outperform the proposed Scoring and Reflective Re-read prompting strategy across the overall evaluation metrics. The improvement in macro precision indicates that ensemble agreement reduces incorrect positive predictions, leading to more reliable stance assignments.
6.4 Generalizability Across Subreddits
To assess generalizability we applied our best-performing method (Mixtral 8x7B + Scoring & Reflective Re-read) to 1,000 unseen comments from three subreddits (567 from r/IsraelPalestine, 233 from r/worldnews, and 200 from r/Palestine) of the original Kaggle dataset. Table 8 compares the predicted stance distributions with the annotated ground truth from the training set.
Table 8: Predicted Stance Distributions on Unseen Data Across Subreddits
Subreddit Stance Annotated (Training) Predicted (Unseen)
r/IsraelPalestine Pro-Israel 49.2% (1,565) 46.0% (261)
Pro-Palestine 30.0% (956) 38.9% (220)
Neutral 20.8% (662) 15.1% (86)
Total 3,183 567
r/worldnews Pro-Israel 50.3% (1,085) 46.0% (107)
Pro-Palestine 15.6% (337) 23.2% (54)
Neutral 34.1% (735) 30.8% (72)
Total 2,157 233
r/Palestine Pro-Israel 6.6% (38) 4.0% (8)
Pro-Palestine 76.0% (438) 91.5% (183)
Neutral 17.4% (100) 4.5% (9)
Total 576 200
The predicted distributions align with the annotated ground truth, with Pro-Israel comments most frequent in r/IsraelPalestine and r/worldnews, and Pro-Palestine comments dominant in r/Palestine. This offers preliminary evidence that the model might be generalizable, though further validation requires more manually annotated data.
7 Broader Impact and Applications
The implications of this research extend beyond model performance and offer valuable insights into the societal dynamics of polarization in conflict zones. Automated detection of ideological stances can support journalists, policymakers, and humanitarian organizations in monitoring public discourse, identifying emerging narratives, and mitigating the spread of misinformation. In the context of the Israel-Palestine conflict, such tools may facilitate balanced reporting and provide early signals of shifts in public sentiment, which can inform dialogue and peace-building efforts. However, it is to be noted that the proposed prompt engineering strategies may not always generalize reliably across different LLMs or datasets. These tools are also not substitutes for qualitative analysis and hence care should be taken during practical application.
At the same time, the deployment of stance detection models raises critical ethical considerations. Misclassification of sensitive content can unintentionally reinforce existing biases or suppress marginalized voices. Additionally, analyzing social media data during ongoing conflicts requires careful attention to privacy and the potential impact on individuals and communities. To ensure responsible use, we recommend integrating human-in-the-loop validation and interdisciplinary oversight when applying these techniques in politically charged contexts.
8 Conclusion
@@
This study presents a comprehensive analysis of ideological stance detection in social media discourse surrounding the Israel–Palestine conflict. Using a manually annotated dataset of Reddit comments, we conducted a comparative evaluation of neural networks, pre-trained language models, and large language models (LLMs) across multiple architectures and prompting configurations. In addition, we explored an LLM ensemble approach based on majority voting to assess whether combining predictions from multiple models improves stance classification performance. The findings highlight the nuanced nature of ideological stance expression in conflict-related discussions and underscore the importance of advanced models and prompt engineering techniques for improving classification performance.
Among the evaluated approaches, Mixtral 8x7B combined with the Scoring and Reflective Re-read prompting strategy demonstrated the best overall performance, achieving improvements in accuracy, recall, and F1-score across stance categories. These results illustrate the limitations of traditional models in capturing subtle and implicit ideological cues in politically sensitive discourse and demonstrate the effectiveness of carefully designed prompting strategies. Our analysis also reveals the complexity of polarization in discussions surrounding the Israel–Palestine conflict, where neutral positions proved particularly challenging to detect (50% precision in baseline models). Furthermore, the stronger performance in detecting Pro-Israel stances compared to Pro-Palestine stances suggests potential asymmetries in how different sides articulate their positions in polarized online environments.
@@
Future work can extend this research by exploring multilingual datasets, incorporating cross-cultural perspectives, investigating fine-tuning approaches for smaller large language models, and examining the ethical implications of ideological stance detection in highly polarized contexts. Overall, the dataset, methodologies, and findings presented in this study contribute to the broader understanding of how natural language processing techniques can be applied to analyze polarization and ideological discourse in conflict-driven social media environments.
Data Availability Statement
All data is available on a Github repository: https://github.com/jami78/Conflict-Bias-Eval.